# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

### 28 Apr 2026 v1

@@ Appendix A Formal Connection to Discrete-time Formulations
In this appendix, we show how the continuous-time stochastic EnKBF formulation (7) and the EnKBS (9) connect to the standard discrete-time Kalman filter and the RTS smoother [32, 33]. By discretizing the continuous-time equations (1), we show that the continuous-time formulations are consistent with the discrete-time filter and smoother up to $O(\tau)$, where $\tau$ is the time increment between observations (assumed uniform, without loss of generality). Starting from the discrete-time filter and smoother, formally taking the limit $\tau\to 0$ yields the continuous-time formulations. As a nonrigorous argument, this is sufficient to highlight the equation structure while avoiding technical details.
Fix a small observation time step $\tau>0$ and define $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\mathbf{x}_{k}=\mathbf{x}(t_{k})$. We consider a first-order semi-implicit Euler–Maruyama type discretization of the nonlinear continuous model (1) [20]:
$\mathbf{x}_{k+1}$ $=\mathbf{x}_{k}+\tau\mathbf{f}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\mathbf{B}_{k},$ (19a)
$\Delta\mathbf{y}_{k+1}$ $=\tau\mathbf{h}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\mathbf{W}_{k+1},$ (19b)
where $\mathbf{B}_{k}$ and $\mathbf{W}_{k+1}$ are independent standard Gaussian vectors, and $\Delta\mathbf{y}_{k+1}=\mathbf{y}_{k+1}-\mathbf{y}_{k}$.
@@
$\mathbf{f}_{k}=\mathbf{F}_{k}\mathbf{x}_{k}+O(\tau),\qquad\mathbf{h}_{k+1}=\mathbf{H}_{k+1}\mathbf{x}_{k+1}+O(\tau).$ (20)
Plugging (20) into (19), we obtain
$\mathbf{x}_{k+1}$ $=(\mathbf{I}+\tau\mathbf{F}_{k})\mathbf{x}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\mathbf{B}_{k}+O(\tau^{2}),$ (21a)
$\Delta\mathbf{y}_{k+1}$ $=\tau\mathbf{H}_{k+1}\mathbf{x}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\mathbf{W}_{k+1}+O(\tau^{2}).$ (21b)
We use the notation $(\cdot)_{k\mid\ell}$ for quantities at time $t_{k}$ conditioned on observations up to time $t_{\ell}$. In this notation, $\mathbf{x}_{k+1\mid k}$ is the prior obtained by dynamical forecast from $t_{k}$ to $t_{k+1}$, and $\mathbf{x}_{k+1\mid k+1}$ is the posterior after assimilating the observation at $t_{k+1}$. Ignoring higher order terms $O(\tau^{2})$, (21) leads to the following ensemble Kalman filter analysis update for $i=1,\dots,m$ [39, 59]:
$\mathbf{x}_{k+1\mid k+1}^{(i)}=\mathbf{x}_{k+1\mid k}^{(i)}+\mathbf{K}_{k+1}\left(\Delta\mathbf{y}_{k+1}-\tau\mathbf{H}_{k+1}\mathbf{x}_{k+1\mid k}^{(i)}-\sqrt{\tau}\mathbf{\Gamma}^{1/2}_{{k+1}}\mathbf{W}^{(i)}_{{k+1}}\right)$, (22)
with $\sqrt{\tau}\mathbf{W}^{(i)}_{{k+1}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ a discrete Brownian increment and $\mathbf{K}_{k+1}$ the ensemble-based Kalman gain:
@@
Here, $\mathbf{P}_{k+1\mid k}$ is the empirical covariance of the ensemble prior $\mathbf{x}_{k+1|k}^{(i)}$. From the analysis step of the Kalman filter, we have [12, 32]
@@
$\mathbf{K}_{k+1}=\mathbf{P}_{k+1\mid k+1}\mathbf{H}_{k+1}^{\top}\mathbf{\Gamma}_{k+1}^{-1}+O(\tau),\qquad\frac{\mathbf{x}_{k+1\mid k+1}^{(i)}-\mathbf{x}_{k+1\mid k}^{(i)}}{\tau}\rightarrow\frac{\,\mathrm{d}\mathbf{x}_{t}^{(i)}}{\,\mathrm{d}s},$
formally taking $\tau\to 0$ in (22) yields the fictitious-time assimilation equation (3). Hence, the stochastic EnKBF formulation (7) for filter-analyzed ensemble members follows by now integrating the derivative-free formulation of (3) (see (5)–(6)) in the forecast dynamics (1a) as an additional drift term [57].
For the smoother, we start from the discrete-time fixed-interval ensemble RTS smoother. Neglecting higher order terms $O(\tau^{2})$, the ensemble-based RTS smoother for (21) reads as [26]
$\mathbf{x}_{k\mid K}^{(i)}=\mathbf{x}_{k\mid k}^{(i)}+\mathbf{C}_{k}\Big(\mathbf{x}_{k+1\mid K}^{(i)}-\mathbf{x}_{k+1\mid k}^{(i)}\Big),\qquad i=1,\dots,m,$ (23)
@@
$\mathbf{C}_{k}:=\mathbf{P}_{k|k}(I+\tau\mathbf{F}_{k}^{\top})\big(\mathbf{P}_{k+1\mid k}\big)^{-1},$ (24)
and the Kalman filter prior covariance by [12, 32]
@@
Hence, its inverse admits the following Neumann series expansion for sufficiently small $\tau$ [64, 79]:
@@
Plugging (25) into (24) and expanding again, we obtain
@@
Denote the smoother correction for the $i$-th ensemble member by $\mathbf{e}_{k}^{(i)}:=\mathbf{x}_{k|K}^{(i)}-\mathbf{x}_{k|k}^{(i)}$. Rearranging (23) gives
$\mathbf{e}_{k+1}^{(i)}=(\mathbf{C}_{k})^{-1}\mathbf{e}_{k}^{(i)}-\left(\mathbf{x}^{(i)}_{k+1\mid k+1}-\mathbf{x}^{(i)}_{k+1\mid k}\right)$. (27)
@@
$\mathbf{x}^{(i)}_{k+1\mid k+1}-\mathbf{x}^{(i)}_{k+1\mid k}\rightarrow\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)},$
where $\mathrm{d}\mathbf{I}^{(i)}:=\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}$ denotes the innovation process, with simulated observation $\mathrm{d}\mathbf{y}^{(i)}$ from (2). Omitting now the subscripts $(\cdot)_{k+1}$ or $(\cdot)_{k+1|k+1}$ in the continuous-time notation, taking the formal limit $\tau\to 0$ and combining (26) and (27) yields [20]
$\mathrm{d}\mathbf{e}^{(i)}=\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}\big)\mathbf{e}^{(i)}\,\mathrm{d}t-\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)}.$ (28)
Here, $\mathbf{e}^{(i)}=\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}$. Similar to the EnKBF, this can also be regarded as an assimilation SDE in a fictitious time, with information from the future state estimates of $\mathbf{x}$ rather than observations from $\mathbf{y}$. Here, the term $\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)}$ counteracts the correction in the forward filter pass. Replacing then the tangent linear drift term $\mathbf{F}\mathbf{x}_{\mathrm{s}}^{(i)}$ with the full dynamics $\mathbf{f}(\mathbf{x}_{\mathrm{s}}^{(i)},t)$ and adding (28) to the filter member equation (7) finally yields the EnKBS dynamics (9):
$\mathrm{d}\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\,\mathrm{d}t.$ (29)
Here, $\mathbf{P}=\mathbf{P}_{\mathrm{f}}$ is the empirical filtering covariance in the continuous-time notation ($\mathbf{P}_{k\mid k}\to\mathbf{P}_{\mathrm{f}},\tau\to 0$). Note that since the smoother dynamics $\mathrm{d}\mathbf{x}_{\mathrm{s}}^{(i)}$ are obtained by directly adding the correction $\mathrm{d}\mathbf{e}^{(i)}$ to the forward filter dynamics $\mathrm{d}\mathbf{x}_{\mathrm{f}}^{(i)}$, the same Wiener process path $\mathbf{B}^{(i)}$ from the filter is naturally preserved in the smoother dynamics 29.
@@ Appendix B Moment Consistency of the EnKBS in the Linear-Gaussian Regime
In this appendix, we demonstrate that the EnKBS equation (9) recovers the classical continuous-time RTS smoother mean and covariance equations for linear-Gaussian systems [33, 56]. We note that all empirical covariance matrices are stochastic objects that depend on the observation path $\mathbf{y}_{[0,T]}$ and the driving noises $\mathbf{B}^{(i)}$, $\mathbf{W}^{(i)}$ for finite ensemble sizes. The analysis below is therefore carried out in a mean-field limit $m\to\infty$, where empirical moments are identical to the true conditional moments [80].
Consider the continuous-time linear-Gaussian setting where the drift and observation maps in (1) are linear: $\mathbf{f}=\mathbf{F}\mathbf{x}$ and $\mathbf{h}=\mathbf{H}\mathbf{x}$. Let $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}:=\mathbf{x}_{\mathrm{s}}^{(i)}-\overline{\mathbf{x}}_{\mathrm{s}}$ and $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}:=\mathbf{x}_{\mathrm{f}}^{(i)}-\overline{\mathbf{x}}_{\mathrm{f}}$ denote the smoother and filter anomalies, respectively. By averaging over ensemble members $i$ in the EnKBS member equation (9) and taking a mean-field (infinite ensemble) limit $m\rightarrow\infty$ [80], we recover the exact mean equation of the RTS smoother:
$\mathrm{d}\overline{\mathbf{x}}_{\mathrm{s}}=\mathbf{F}(t)\overline{\mathbf{x}}_{\mathrm{s}}\,\mathrm{d}t+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\overline{\mathbf{x}}_{\mathrm{s}}-\overline{\mathbf{x}}_{\mathrm{f}}\big)\,\mathrm{d}t.$ (30)
Subtracting (30) from the EnKBS member equation (9) yields the anomaly dynamics:
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}=\left((\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}})\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\right)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}.$ (31)
We now define the smoother posterior covariance as $\mathbf{P}_{\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$ and the filter–smoother cross-covariance as $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$. Applying Itô’s product rule to $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$ yields [51, 50]:
$\mathrm{d}\big(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big)=(\,\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\,\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\mathbf{\Sigma}\,\mathrm{d}t,$ (32)
where $\mathbf{\Sigma}\,\mathrm{d}t=\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})^{\top}\big]$ emerges as the Itô correction term. Taking the expectation gives the continuous-time covariance equation:
@@
Therefore, the standard continuous-time RTS covariance equation (33) is recovered if

### 03 May 2026 v2

@@ Appendix A Formal Connection to Discrete-time Formulations
In this appendix, we show how the continuous-time stochastic EnKBF formulation Eq. (7) and the EnKBS Eq. (9) connect to the standard discrete-time Kalman filter and the RTS smoother (Kalman, 1960; Rauch et al., 1965). By discretizing the continuous-time equations Eq. (1), we show that the continuous-time formulations are consistent with the discrete-time filter and smoother up to $O(\tau)$, where $\tau$ is the time increment between observations (assumed uniform, without loss of generality). Starting from the discrete-time filter and smoother, formally taking the limit $\tau\to 0$ yields the continuous-time formulations. As a nonrigorous argument, this is sufficient to highlight the equation structure while avoiding technical details.
Fix a small observation time step $\tau>0$ and define $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\vec{x}_{k}=\vec{x}(t_{k})$. We consider a first-order semi-implicit Euler–Maruyama type discretization of the nonlinear continuous model Eq. (1) (Law et al., 2015):
$\vec{x}_{k+1}$ $=\vec{x}_{k}+\tau\vec{f}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\vec{B}_{k},$ (19a)
$\Delta\vec{y}_{k+1}$ $=\tau\vec{h}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\vec{W}_{k+1},$ (19b)
where $\vec{B}_{k}$ and $\vec{W}_{k+1}$ are independent standard Gaussian vectors, and $\Delta\vec{y}_{k+1}=\vec{y}_{k+1}-\vec{y}_{k}$.
@@
$\vec{f}_{k}=\mathbf{F}_{k}\vec{x}_{k}+O(\tau),\qquad\vec{h}_{k+1}=\mathbf{H}_{k+1}\vec{x}_{k+1}+O(\tau).$ (20)
Plugging Eq. (20) into Eq. (19), we obtain
$\vec{x}_{k+1}$ $=(\mathbf{I}+\tau\mathbf{F}_{k})\vec{x}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\vec{B}_{k}+O(\tau^{2}),$ (21a)
$\Delta\vec{y}_{k+1}$ $=\tau\mathbf{H}_{k+1}\vec{x}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\vec{W}_{k+1}+O(\tau^{2}).$ (21b)
We use the notation $(\cdot)_{k\mid\ell}$ for quantities at time $t_{k}$ conditioned on observations up to time $t_{\ell}$. In this notation, $\vec{x}_{k+1\mid k}$ is the prior obtained by dynamical forecast from $t_{k}$ to $t_{k+1}$, and $\vec{x}_{k+1\mid k+1}$ is the posterior after assimilating the observation at $t_{k+1}$. Ignoring higher order terms $O(\tau^{2})$, Eq. (21) leads to the following ensemble Kalman filter analysis update for $i=1,\dots,m$ (Burgers et al., 1998; Reich, 2011):
$\vec{x}_{k+1\mid k+1}^{(i)}=\vec{x}_{k+1\mid k}^{(i)}+\mathbf{K}_{k+1}\left(\Delta\vec{y}_{k+1}-\tau\mathbf{H}_{k+1}\vec{x}_{k+1\mid k}^{(i)}-\sqrt{\tau}\mathbf{\Gamma}^{1/2}_{{k+1}}\vec{W}^{(i)}_{{k+1}}\right)$, (22)
with $\sqrt{\tau}\vec{W}^{(i)}_{{k+1}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ a discrete Brownian increment and $\mathbf{K}_{k+1}$ the ensemble-based Kalman gain:
@@
Here, $\mathbf{P}_{k+1\mid k}$ is the empirical covariance of the ensemble prior $\vec{x}_{k+1|k}^{(i)}$. From the analysis step of the Kalman filter, we have (Asch et al., 2016; Kalman, 1960)
@@
$\mathbf{K}_{k+1}=\mathbf{P}_{k+1\mid k+1}\mathbf{H}_{k+1}^{\top}\mathbf{\Gamma}_{k+1}^{-1}+O(\tau),\qquad\frac{\vec{x}_{k+1\mid k+1}^{(i)}-\vec{x}_{k+1\mid k}^{(i)}}{\tau}\rightarrow\frac{\mathrm{d}\vec{x}_{t}^{(i)}}{\mathrm{d}s},$
formally taking $\tau\to 0$ in Eq. (22) yields the fictitious-time assimilation equation Eq. (3). Hence, the stochastic EnKBF formulation Eq. (7) for filter-analyzed ensemble members follows by now integrating the derivative-free formulation of Eq. (3) (see Eq. (5)–Eq. (6)) in the forecast dynamics Eq. (1a) as an additional drift term (Bergemann and Reich, 2012).
For the smoother, we start from the discrete-time fixed-interval ensemble RTS smoother. Neglecting higher order terms $O(\tau^{2})$, the ensemble-based RTS smoother for Eq. (21) reads as (Evensen and Van Leeuwen, 2000)
$\vec{x}_{k\mid K}^{(i)}=\vec{x}_{k\mid k}^{(i)}+\mathbf{C}_{k}\Big(\vec{x}_{k+1\mid K}^{(i)}-\vec{x}_{k+1\mid k}^{(i)}\Big),\qquad i=1,\dots,m,$ (23)
@@
$\mathbf{C}_{k}:=\mathbf{P}_{k|k}(\mathbf{I}+\tau\mathbf{F}_{k}^{\top})\big(\mathbf{P}_{k+1\mid k}\big)^{-1},$ (24)
and the Kalman filter prior covariance by (Asch et al., 2016; Kalman, 1960)
@@
Hence, its inverse admits the following Neumann series expansion for sufficiently small $\tau$ (Horn and Johnson, 1985; Folland, 1999):
@@
Plugging Eq. (25) into Eq. (24) and expanding again, we obtain
@@
Denote the smoother correction for the $i$-th ensemble member by $\vec{e}_{k}^{(i)}:=\vec{x}_{k|K}^{(i)}-\vec{x}_{k|k}^{(i)}$. Rearranging Eq. (23) gives
$\vec{e}_{k+1}^{(i)}=(\mathbf{C}_{k})^{-1}\vec{e}_{k}^{(i)}-\left(\vec{x}^{(i)}_{k+1\mid k+1}-\vec{x}^{(i)}_{k+1\mid k}\right)$. (27)
@@
$\vec{x}^{(i)}_{k+1\mid k+1}-\vec{x}^{(i)}_{k+1\mid k}\rightarrow\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathrm{d}\vec{I}^{(i)},$
where $\mathrm{d}\vec{I}^{(i)}:=\mathrm{d}\vec{y}-\mathrm{d}\vec{y}^{(i)}$ denotes the innovation process, with simulated observation $\mathrm{d}\vec{y}^{(i)}$ from Eq. (2). Omitting now the subscripts $(\cdot)_{k+1}$ or $(\cdot)_{k+1|k+1}$ in the continuous-time notation, taking the formal limit $\tau\to 0$ and combining Eq. (26) and Eq. (27) yields (Law et al., 2015)
$\mathrm{d}\vec{e}^{(i)}=\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}\big)\vec{e}^{(i)}\mathrm{d}t-\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathrm{d}\vec{I}^{(i)}.$ (28)
Here, $\vec{e}^{(i)}=\vec{x}_{\mathrm{s}}^{(i)}-\vec{x}_{\mathrm{f}}^{(i)}$. Similar to the EnKBF, this can also be regarded as an assimilation SDE in a fictitious time, with information from the future state estimates of $\vec{x}$ rather than observations from $\vec{y}$. Here, the term $\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathrm{d}\vec{I}^{(i)}$ counteracts the correction in the forward filter pass. Replacing then the tangent linear drift term $\mathbf{F}\vec{x}_{\mathrm{s}}^{(i)}$ with the full dynamics $\vec{f}(\vec{x}_{\mathrm{s}}^{(i)},t)$ and adding Eq. (28) to the filter member equation Eq. (7) finally yields the EnKBS dynamics Eq. (9):
$\mathrm{d}\vec{x}_{\mathrm{s}}^{(i)}=\vec{f}\big(\vec{x}_{\mathrm{s}}^{(i)},t\big)\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\vec{x}_{\mathrm{s}}^{(i)}-\vec{x}_{\mathrm{f}}^{(i)}\big)\mathrm{d}t.$ (29)
Here, $\mathbf{P}=\mathbf{P}_{\mathrm{f}}$ is the empirical filtering covariance in the continuous-time notation ($\mathbf{P}_{k\mid k}\to\mathbf{P}_{\mathrm{f}},\tau\to 0$). Note that since the smoother dynamics $\mathrm{d}\vec{x}_{\mathrm{s}}^{(i)}$ are obtained by directly adding the correction $\mathrm{d}\vec{e}^{(i)}$ to the forward filter dynamics $\mathrm{d}\vec{x}_{\mathrm{f}}^{(i)}$, the same Wiener process path $\vec{B}^{(i)}$ from the filter is naturally preserved in the smoother dynamics Eq. (29).
@@ Appendix B Moment Consistency of the EnKBS in the Linear-Gaussian Regime
In this appendix, we demonstrate that the EnKBS equation Eq. (9) recovers the classical continuous-time RTS smoother mean and covariance equations for linear-Gaussian systems (Rauch et al., 1965; Särkkä and Svensson, 2023). We note that all empirical covariance matrices are stochastic objects that depend on the observation path $\vec{y}_{[0,T]}$ and the driving noises $\vec{B}^{(i)}$, $\vec{W}^{(i)}$ for finite ensemble sizes. The analysis below is therefore carried out in a mean-field limit $m\to\infty$, where empirical moments are identical to the true conditional moments (Calvello et al., 2025).
Consider the continuous-time linear-Gaussian setting where the drift and observation maps in Eq. (1) are linear: $\vec{f}=\mathbf{F}\vec{x}$ and $\vec{h}=\mathbf{H}\vec{x}$. Let $\vec{\hat{x}}_{\mathrm{s}}^{(i)}:=\vec{x}_{\mathrm{s}}^{(i)}-\overline{\vec{x}}_{\mathrm{s}}$ and $\vec{\hat{x}}_{\mathrm{f}}^{(i)}:=\vec{x}_{\mathrm{f}}^{(i)}-\overline{\vec{x}}_{\mathrm{f}}$ denote the smoother and filter anomalies, respectively. By averaging over ensemble members $i$ in the EnKBS member equation Eq. (9) and taking a mean-field (infinite ensemble) limit $m\rightarrow\infty$ (Calvello et al., 2025), we recover the exact mean equation of the RTS smoother:
$\mathrm{d}\overline{\vec{x}}_{\mathrm{s}}=\mathbf{F}\overline{\vec{x}}_{\mathrm{s}}\mathrm{d}t+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\overline{\vec{x}}_{\mathrm{s}}-\overline{\vec{x}}_{\mathrm{f}}\big)\mathrm{d}t.$ (30)
Subtracting Eq. (30) from the EnKBS member equation Eq. (9) yields the anomaly dynamics:
$\mathrm{d}\vec{\hat{x}}_{\mathrm{s}}^{(i)}=\left((\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}})\vec{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\vec{\hat{x}}_{\mathrm{f}}^{(i)}\right)\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}.$ (31)
We now define the smoother posterior covariance as $\mathbf{P}_{\mathrm{s}}=\mathbb{E}\big[\vec{\hat{x}}_{\mathrm{s}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$ and the filter–smoother cross-covariance as $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\vec{\hat{x}}_{\mathrm{f}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$. Applying Itô’s product rule to $\vec{\hat{x}}_{\mathrm{s}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$ yields (Øksendal, 2003; Särkkä and Solin, 2019):
$\mathrm{d}\big(\vec{\hat{x}}_{\mathrm{s}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big)=(\mathrm{d}\vec{\hat{x}}_{\mathrm{s}}^{(i)})(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\vec{\hat{x}}_{\mathrm{s}}^{(i)}(\mathrm{d}\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\mathbf{\Sigma}\mathrm{d}t,$ (32)
where $\mathbf{\Sigma}\mathrm{d}t=\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)})(\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)})^{\top}\big]$ emerges as the Itô correction term. Taking the expectation gives the continuous-time covariance equation:
@@
Therefore, the standard continuous-time RTS covariance equation Eq. (33) is recovered if